OpenStudy (anonymous):

Find all vectors v such that (1,2,1) x v = (3,1,-5) (x=cross product)

OpenStudy (amistre64):

teh cross product is used to find the normal of a plane; therefore all the vectors the reside in the plane will cross to get the desired results

OpenStudy (amistre64):

the equation of the plane is the set of all vectors that can be crossed :)

OpenStudy (amistre64):

if we cross these 2 that are given, we will find at least one of the vectors that reside in the plane

OpenStudy (anonymous):

would the answer be the plane 3(x-1)+1(y-2)-5(z-1)?

OpenStudy (amistre64):

id say yes; since any 2 vectors on the place cross to get the normal

OpenStudy (amistre64):

wait, we havnt actually got any points to define a plane by tho; so this needs a little more thoughts

OpenStudy (anonymous):

ok, well v has to be orthogonal to (3,1,-5)

OpenStudy (amistre64):

(1,2,1) (3,1,-5) ------- x= -11 y = 8 z = -4 yes, i agree with that

OpenStudy (anonymous):

you did the cross product of the two?

OpenStudy (amistre64):

yeah, that gives me a vector that is at least in the same plane as the other :)

OpenStudy (amistre64):



OpenStudy (amistre64):

these (...) are vectors right? i mean im just assuming that they are. im used to a vector being denoted as <...>

OpenStudy (anonymous):

could it be the only possible vector? yes they are vectors

OpenStudy (amistre64):

no, it cant be, we can add this new one to the first and cross it to get the normal just as well

OpenStudy (amistre64):

( 1, 2 ,1 ) (-11,8,-4) ---------- <-10,10,-3> = (1/10)<-1,1,-.3> would you agree?

OpenStudy (anonymous):

hmm, i see <0, -7, 30>, cross product right?

OpenStudy (anonymous):

oh sorry, I see, yes

OpenStudy (anonymous):

why did u add them?

OpenStudy (amistre64):

1 ,2 ,1 -10,10,-3 ------------- x = -16 y = -7 z = 30 <-16,-7, 30> <16, 7, -30> /6 /6 /6 ----------- hmmmm <3,1,-5> am i dong the math right is what i gottta wonder now

OpenStudy (amistre64):

i add them to find another vector that i know is in the plane to use as a test to cross with

OpenStudy (amistre64):

the problem with vectors is that they can exist anyplace in the plane; they move about

OpenStudy (amistre64):



OpenStudy (anonymous):

if we suppose the first two vectors are in the xy plane, then the 2nd vector can be any non zero vector other than the 1st vector right?

OpenStudy (amistre64):

sure, i dont see why not at the moment

OpenStudy (anonymous):

I think this is easier than we think, perhaps you just do cross product of <1,2,1> and v and make it equal to <3,1,-5>

OpenStudy (anonymous):

then do matrix

OpenStudy (amistre64):

(1,2,1) (3,1,-5) ------- x= -10 - 1 = -11 y = 3 +5 = 8 z = 1- 6 = -5 that might be better

OpenStudy (amistre64):

my z last time was 1 - (3+2) instead of 1 - 3(2) lol

OpenStudy (amistre64):

so v can = t<11,-8,5>

OpenStudy (anonymous):

ok cool thanks

OpenStudy (amistre64):

but all the vectors? as far as i can see it, there are infinite number of them all in planes that are parallel to each other

OpenStudy (anonymous):

ok thatt's a possible answer, the next part asks why there is no vector v such that =<3,1,5>

OpenStudy (amistre64):

(1,2,1) (a, b, c) ------- x = 2c -b = 3 y = a - c = 1 z = b - 2a = -5 how about them apples?

OpenStudy (anonymous):

niccceee

OpenStudy (amistre64):

at the moment thats the best I got :)

OpenStudy (anonymous):

thanks, Ive got to head to bed tho, have a good one

OpenStudy (amistre64):

and i dont think 3,1,5 is possible since it aint gonna be orthogonal to the 3,1,-5